What is 25% of 62.00
Answer by Maths68(1474)

25% of 62.00 = (25/100) * 62
25% of 62.00 = (1/4) * 62
25% of 62.00 = 62/4
25% of 62.00 = 31/2
25% of 62.00 = 15.5